OpenStudy (anonymous):

Integrate [64/(16+xˆ2)]-1 dx, lower limit of -2√2 and upper limit 2√2. Should I use U substitution, or another method?

OpenStudy (sirm3d):

use the standard formula \[\Large \int \frac{\mathrm dx}{a^2+x^2}=\frac{1}{a} \tan^{-1}\frac{x}{a}\]